OpenStudy (anonymous):

Find the fourth roots of 256.

OpenStudy (karatechopper):

wat is that? is tht 4 times the sqrt of 156?

OpenStudy (anonymous):

number that when multiplied by itself four times (raised to the power 4) equals 256.

OpenStudy (karatechopper):

o!

OpenStudy (anonymous):

can you solve the equation:\[x^4=256\]? thats pretty much what the question is asking

OpenStudy (anonymous):

i got 4 , -4

OpenStudy (anonymous):

Thats correct :) those are the real fourth roots of 256, since:\[4^4=256,(-4)^4=256\]